# A coisotropic embedding theorem for pre-multisymplectic manifolds

### v1

@@ 3.1 Statement
Let $(\mathcal{M},\omega)$ be a $k$-pre-multisymplectic manifold. Then, there exists a multisymplectic manifold $\tilde{\mathcal{M}}$, referred to as the multisymplectic thickening of $\mathcal{M}$, and an embedding $\mathfrak{i}\;:\;\;\mathcal{M}\hookrightarrow\tilde{\mathcal{M}}$, such that:
@@
    $\tilde{\mathcal{M}}$ is equipped with a $k$-multisymplectic form $\tilde{\omega}$.
@@
    $(\mathcal{M},\omega)$ becomes a $(k-1)$-coisotropic submanifold of (~$(\tilde{\mathcal{M}},\tilde{\omega})$. In particular $\mathfrak{i}^{\star}\tilde{\omega}\,=\,\omega$.
3.2 Proof
Consider a $d$-dimensional $k$-pre-multisymplectic manifold $(\mathcal{M},\,\omega)$. In the system of coordinates adapted to the foliation induced by the kernel of $\omega$ described in Section 2, $\omega$ reads:
@@ 3.2 Proof
Consider the bundle ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ described in Section 2, with the system of local coordinates
$\left\{\,q^{j},\,p_{I^{A}_{i}}\,\right\}_{j=1,...,d;\,I^{A}_{i}\in\mathscr{I}}\,.$ (34)
Consider the $(k-1)$-form on ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ defined by
${\Theta_{0}^{(k-1)}}_{p}(V_{1},...,V_{k-1})\,=\,p(\pi_{\star}V_{1},...,\pi_{\star}V_{k-1})\,,$ (35)
where $p$ has to be understood as a point in ${\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$ on the left hand side and as a (transversal) $(k-1)$-form on $\mathcal{M}$ on the right hand side, $\pi$ is the projection from ${\Lambda^{k}}^{\perp}_{R}(\mathcal{M})$ to $\mathcal{M}$, and $V_{j}\in\mathbf{T}_{p}{\Lambda^{k}}^{\perp}_{R}(\mathcal{M})\,,\;\;\;\forall\,\,j=1,...,k-1$. In the system of local coordinates chosen
${\Theta_{0}^{(k-1)}}\,=\,\sum_{I^{A}_{i}}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge{\rm d}\rho^{j_{k-1}}\,.$ (36)
@@
$\begin{split}\tilde{\omega}\,:=\,\pi^{\star}\omega+{\rm d}\Theta^{(k-1)}_{0}\,.\end{split}$ (37)
In the system of local coordinates chosen
$\begin{split}\tilde{\omega}\,=\,&\omega_{a_{1}...a_{k}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k}}+\\ +&\sum_{I^{A}_{i}}{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+\\ +&\sum_{I^{A}_{i},\,j}(-1)^{j+1}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,.\end{split}$ (38)
@@
To prove whether it is also non-degenerate, we will consider its contraction along the following basis of vector fields on $\tilde{\mathcal{M}}$:
$\left\{\,\frac{\partial}{\partial f^{A}},\,W_{a},\,\frac{\partial}{\partial p^{I^{A}_{i}}}\,\right\}_{a=1,...,l;\,A=1,...,r;\,I^{A}_{i}\in\mathscr{I}}\,.$ (39)
Via a direct computation one shows
$i_{\frac{\partial}{\partial f^{A}}}\tilde{\omega}\,=\,$ $\sum_{I^{A}_{i}}(-1)^{i}{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\underbrace{\widehat{\theta^{A}}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$ (40)
$+$ $\sum_{I^{A}_{i},\,j}(-1)^{j+i}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\widehat{\theta^{A}}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$
-- $\sum_{I^{A}_{i},\,j}p_{I^{A}_{i}}\,{\rm d}_{V}\theta^{B}\left(\,\frac{\partial}{\partial f^{A}}\,\right)\,\rho^{j_{1}}\wedge...\wedge\underbrace{\widehat{{\rm d}\theta^{B}}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,,$
$i_{W_{a}}\tilde{\omega}\,=\,$ $\omega_{a_{1}...a...a_{k}}\,{\rm d}x^{a_{1}}\wedge...\wedge\widehat{{\rm d}x^{a}}\wedge...\wedge{\rm d}x^{a_{k}}+$ (41)
$+$ $\sum_{I^{A}_{i}}(-1)^{j}{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\underbrace{{\rm d}x^{a}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$
$+$ $\sum_{I^{A}_{i},\,j}(-1)^{i+l}p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\widehat{{\rm d}x^{l}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+$
-- $\sum_{I^{A}_{i},\,j}{\rm d}_{H}\theta^{B}(W^{a})\rho^{j_{1}}\wedge...\wedge\underbrace{\widehat{{\rm d}\theta^{B}}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,,$
$i_{\frac{\partial}{\partial p_{I^{A}_{i}}}}\tilde{\omega}\,=\,$ $\rho^{j_{1}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,,$ (42)
where ${\rm d}_{V}\theta^{A}\,=\,{\rm d}\theta^{A}(P(\,\cdot\,),\,P(\,\cdot\,))$ and ${\rm d}_{H}\theta^{A}\,=\,{\rm d}\theta^{A}(R(\,\cdot\,),\,R(\,\cdot\,))$. Since none of the right-hand sides of the equations above vanish, one concludes that $\tilde{\omega}$ is non-degenerate.
@@
Note that the departing $k$-pre-multisymplectic manifold $\mathcal{M}$ can be embedded into $\Lambda^{k-1}(\mathcal{M})$ as the zero-section $\sigma_{0}\,=:\,\mathfrak{i}$ of the projection $\pi$. A straightforward computation also shows that
$\mathfrak{i}^{\star}\tilde{\omega}\,=\,\omega\,.$ (43)
To show that $\mathcal{M}$ is coisotropic inside $\tilde{\mathcal{M}}$, let us consider a vector field tangent to $\mathcal{M}$
$W\,=\,W^{j}(q)\frac{\partial}{\partial q^{j}}\,,$ (44)
and a vector field tangent to $\tilde{\mathcal{M}}$
$V\,=\,V^{j}(q,\,p)\frac{\partial}{\partial q^{j}}+V^{I^{A}_{i}}(q,\,p)\frac{\partial}{\partial p_{I^{A}_{i}}}\,,$ (45)
and let us compute
@@
A direct computation shows that
$\begin{split}i_{V\wedge W}\tilde{\omega}\,=\,&\omega_{a_{1}...a_{s}...a_{l}...a_{k}}\,(-1)^{j+k+1}\,V^{s}W^{l}\,{\rm d}x^{a_{1}}\wedge...\wedge\widehat{{\rm d}x^{a_{s}}}\wedge...\wedge\widehat{{\rm d}x^{a_{l}}}\wedge...\wedge{\rm d}x^{a_{k}}+\\ +&\sum_{I^{A}_{i}}(-1)^{l+1}V^{I^{A}_{i}}W^{l}\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+\\ +&\sum_{I^{A}_{i}}(-1)^{s+l}V^{s}W^{l}\,{\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{s}}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}+\\ +&\sum_{I^{A}_{i},\,j}(-1)^{j+1+l+s}V^{s}W^{l}\,p_{I^{A}_{i}}\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{s}}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{{\rm d}\theta^{B}}_{j}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}\,.\end{split}$ (47)
The term involving
${\rm d}p_{I^{A}_{i}}\wedge\rho^{j_{1}}\wedge...\wedge\widehat{\rho^{j_{s}}}\wedge...\wedge\widehat{\rho^{j_{l}}}\wedge...\wedge\underbrace{\theta^{A}}_{i}\wedge...\wedge\rho^{j_{k-1}}$ (48)
is independent on the others. Thus, imposing $i_{V\wedge W}\tilde{\omega}$ to vanish for all $W\in\mathscr{X}(\mathcal{M})$ forces all the $V^{j}$ to be zero. Imposing this condition into the other terms leads to the fact that also all the $V^{I^{A}_{i}}$ vanish. Consequently, $\mathbf{T}\mathcal{M}^{\perp,1}$ is empty, and, thus, $\mathcal{M}$ is trivially $1$-coisotropic inside $\tilde{\mathcal{M}}$.
The arguments above can be iterated up to the order $(k-1)$ to show that $\mathcal{M}$ is $(k-1)$-coisotropic inside $\tilde{\mathcal{M}}$.

### v2

@@ 2 Coisotropic embeddings of pre-multisymplectic manifolds
Let $(\mathcal{M},\omega)$ be a pre-$k$-plectic manifold, with $k>2$. Assume $K\,=\,\mathrm{ker}\,\omega$ to have constant rank. Then, there exists a multisymplectic manifold $\tilde{\mathcal{M}}$, referred to as the multisymplectic thickening of $\mathcal{M}$, and an embedding $\mathfrak{i}\;:\;\;\mathcal{M}\hookrightarrow\tilde{\mathcal{M}}$, such that:
@@
    $\tilde{\mathcal{M}}$ is equipped with a multisymplectic $k$-form $\tilde{\omega}$.
@@
[…]
Proof.
Consider a $d$-dimensional pre-$k$-plectic manifold $(\mathcal{M},\,\omega)$. In the system of coordinates adapted to the foliation induced by the kernel of $\omega$ described in Section˜1, $\omega$ reads:
@@
Consider the bundle ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ described in Section˜1, with the system of local coordinates
$\left\{\,q^{j},\,p_{A_{1}...A_{k}},\,p_{a_{1}A_{2}...A_{k}},\,...,\,p_{a_{1}...a_{k-1}A_{k}}\,\right\}_{j=1,...,d;a_{k}=1,...,r;A_{k}=1,...,l}\,.$ (35)
Let us also denote by $\tau:{\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})\to\mathcal{M}$ the canonical bundle projection. Consider the (k1)(k-1)-form on ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ defined by
${\Theta_{0}^{(k-1)}}_{p}(V_{1},...,V_{k-1})\,=\,p(\tau_{\star}V_{1},...,\tau_{\star}V_{k-1})\,$, (36)
where $p$ has to be understood as a point in ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ on the left hand side and as a (transversal) (k1)(k-1)-covector on $\mathcal{M}$ on the right hand side, $\tau_{\star}$ is the push-forward of $\tau$, and $V_{j}\in\mathbf{T}_{p}{\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})\,,\;\;\;\forall\,\,j=1,...,k-1$. In the system of local coordinates chosen
$\begin{split}{\Theta_{0}^{(k-1)}}\,=\,&p_{A_{1}\,...\,A_{k-1}}\theta^{A_{1}}\wedge...\wedge\theta^{A_{k-1}}+\\ &p_{a_{1}\,A_{2}\,...\,A_{k-1}}{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge...\wedge\theta^{A_{k-1}}+\\ &\;\;\vdots\\ &p_{a_{1}\,...\,a_{k-2}\,A_{k-1}}{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k-2}}\wedge\theta^{A_{k-1}}\,.\end{split}$ (37)
@@
[…]
@@
To prove that it is also non-degenerate, we will consider its contraction along a vector field $X$ written in the basis
$\left\{\,H_{a}=\frac{\partial}{\partial x^{a}}+{P_{x}}^{A}_{a}\frac{\partial}{\partial f^{A}},\,V^{A}=\frac{\partial}{\partial f^{A}},\,\frac{\partial}{\partial p_{A_{1}\,...\,A_{k-1}}},\,\frac{\partial}{\partial p_{a_{1}\,A_{2}\,...\,A_{k-1}}},\,...\,,\,\frac{\partial}{\partial p_{a_{1}\,...\,a_{k-2}\,A_{k-1}}}\,\right\}\,$,
namely
[…]
and prove that
$i_{X}\tilde{\omega}\,=\,0\;\;\forall\,\,X\in\mathfrak{X}({\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M}))\;\;\implies\;\;X=0$
The contraction $i_{X}\widetilde{\omega}$ reads
$\begin{split}i_{X}\widetilde{\omega}\,=\,&{C^{0}}_{A_{1}\,...\,A_{k-1}}\,\theta^{A_{1}}\wedge...\wedge\theta^{A_{k-1}}+\\ &{C^{1}}_{a_{1}\,A_{2}\,...\,A_{k-1}}\,{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge\,...\,\wedge\theta^{A_{k-1}}+\\ &\;\;\vdots\\ &{C^{k-1}}_{a_{1}\,...\,a_{k-1}}\,{\rm d}x^{a_{1}}\wedge\,...\,{\rm d}x^{a_{k-1}}\,+\\ &{D^{0}}^{A_{1}}\,{\rm d}p_{A_{1}\,...\,A_{k-1}}\wedge\theta^{A_{2}}\wedge\,...\,\wedge\theta^{A_{k-1}}+\\ &{D^{1}_{0}}^{a_{1}}\,{\rm d}p_{a_{1}\,A_{2}\,...\,A_{k-1}}\wedge\theta^{A_{2}}\wedge\,...\,\wedge\theta^{A_{k-1}}+\\ &{D^{1}_{1}}^{A_{2}}{\rm d}p_{a_{1}\,A_{2}\,...\,A_{k-1}}\wedge{\rm d}x^{a_{1}}\wedge\theta^{A_{3}}\wedge\,...\,\wedge\theta^{A_{k-1}}+\\ &\;\;\vdots\\ &{D^{k-1}_{k-2}}^{a_{1}}{\rm d}p_{a_{1}\,...\,a_{k-2}A_{k-1}}\wedge{\rm d}x^{a_{2}}\wedge...\wedge{\rm d}x^{a_{k-2}}\wedge\theta^{A_{k-1}}+\\ &{D^{k-1}_{k-1}}^{A_{k-1}}{\rm d}p_{a_{1}\,...\,a_{k-2}A_{k-1}}\wedge{\rm d}x^{a_{1}}\wedge...\wedge{\rm d}x^{a_{k-2}}\,.\end{split}$ (40)
In particular, ${D^{0}}^{A_{1}}$ reads
${D^{0}}^{A_{1}}\,=\,-X^{A_{1}}\,$, (41)
and, since
${\rm d}p_{A_{1}\,...\,A_{k-1}}\wedge\theta^{A_{2}}\wedge\,...\,\wedge\theta^{A_{k-1}}$ (42)
is independent of all the other forms of the above decomposition, the condition $i_{X}\widetilde{\omega}\,=\,0\,,\;\;\;\forall\,\,X$ gives
$X^{A}\,=\,0\,,\;\;\forall\,\,A=1,...,l\,.$ (43)
The latter condition implies
${D^{1}_{1}}^{A_{2}}\,=\,-X^{A_{2}}\,=\,0\,$, (44)
${D^{1}_{0}}^{a_{1}}\,=\,-X^{a_{1}}\,.$ (45)
Again, given that
${\rm d}p_{a_{1}\,A_{2}\,...\,A_{k-2}}\wedge\theta^{A_{2}}\wedge\,...\,\wedge\theta^{A_{k-1}}$ (46)
is independent of all the other forms appearing in (40), the condition $i_{X}\widetilde{\omega}\,=\,0\,,\;\;\;\forall\,\,X$ implies
$X^{a}\,=\,0\,,\;\;\forall\,\,a=1,...,r\,.$ (47)
Conditions (43) and (47) imply in turn that all the $D^{j}_{k}$} vanish and that
${C^{0}}_{A_{1}\,...\,A_{k-1}}\,$ $=\,X_{A_{1}\,...\,A_{k-1}}\,$, (48)
${C^{1}}_{a_{1}\,A_{2}\,...\,A_{k-1}}\,$ $=\,X_{a_{1}\,A_{2}\,...\,A_{k-1}}\,$, (49)
$\vdots$ (50)
${C^{k-2}}_{a_{1}\,...\,a_{k-2}\,A_{k-1}}\,$ $=\,X_{a_{1}\,...\,a_{k-2}\,A_{k-1}}\,$, (51)
${C^{k-1}}_{a_{1}\,...\,a_{k-1}}\,$ = $=\,0\,.$ (52)
Since the $C^{j}_{k}$ are coefficients of independent (k1)(k-1)-forms, the condition $i_{X}\widetilde{\omega}\,=\,0\,,\;\;\forall\,\,X$ implies
$X_{A_{1}\,...\,A_{k-1}}\,$ = $=\,0\,$, (53)
$X_{a_{1}\,A_{2}\,...\,A_{k-1}}\,$ = $=\,0\,$, (54)
$\vdots$
$X_{a_{1}\,...\,a_{k-2}\,A_{k-1}}\,$ = $=\,0\,.$ (55)
The chain of conditions (43), (47), (53), (54), …, (55) prove that
$i_{X}\widetilde{\omega}\,=\,0\,,\;\;\forall\,\,X\in\mathfrak{X}({\Lambda^{k-1}}^{\perp}_{R}(M))\;\;\implies\;\;X\,=\,0\,$, (56)
namely, that $\widetilde{\omega}$ is non-degenerate, and, thus, multisymplectic.
Thus, the multisymplectic manifold we were searching for is $\tilde{\mathcal{M}}\,=\,{\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$ and $\mathfrak{i}$ is the zero-section of $\tau$.
A straightforward computation also shows that
$\mathfrak{i}^{\star}\tilde{\omega}\,=\,\omega\,.$ (57)
To prove that $\mathcal{M}$ is a (k1)(k-1)-coisotropic submanifold of ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$, having in mind Definition˜1.4, consider the contraction $\tilde{\omega}_{m}(X,\,W_{1},...,\,W_{k-1})$ for $m\in\mathcal{M}$, and $W_{1},\,W_{2},\,...,\,W_{k-1}\in\mathbf{T}_{m}\mathcal{M}$
$W_{j}\,=\,{W_{j}}^{a}H_{a}\bigr|_{m}+{W_{j}}^{A}\frac{\partial}{\partial f^{A}}\Bigr|_{m}\,.$ (58)
The contraction above is
$\begin{split}\tilde{\omega}_{m}(X,\,W_{1},...,\,W_{k-1})\,=\,&\tau^{*}\omega_{m}(X,\,W_{1},...,\,W_{k-1})\,+\\ &X_{A_{1}\,...\,A_{k-1}}\left(\,\theta^{A_{1}}\wedge...\wedge\theta^{A_{k-1}}\,\right)(W_{1},...,\,W_{k-1})\,+\\ &X_{a_{1}\,A_{2}\,...\,A_{k-1}}\left(\,{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge...\wedge\theta^{A_{k-1}}\,\right)(\,W_{1},...,\,W_{k-1})\,+\\ &\;\;\vdots\\ &X_{a_{1}\,...\,a_{k-2}\,A_{k-1}}\left(\,{\rm d}x^{a_{1}}\wedge...{\rm d}x^{a_{k-2}}\wedge\theta^{A_{k-1}}\,\right)(W_{1},...,\,W_{k-1})\,.\end{split}$ (59)
First, consider all the $W_{j}$ to be vertical, that is, obeying ${W_{j}}^{a}\,=\,0$. The contraction above reads
$X_{A_{1}\,...\,A_{k-1}}\left(\,\theta^{A_{1}}\wedge...\wedge\theta^{A_{k-1}}\,\right)(W_{1},...,\,W_{k-1})\,.$ (60)
The condition that it has to vanish for all the $W_{j}$ gives
$X_{A_{1}\,...\,A_{k-1}}\,=\,0\,.$ (61)
Let now $W_{1}$ be horizontal, i.e., $W_{1}\,=\,{W_{1}}^{a}H_{a}$, and the remaining $W_{j}$ be vertical as in the previous case. Here, the contraction reads
$X_{a_{1}\,A_{2}\,...\,A_{k-1}}\left(\,{\rm d}x^{a_{1}}\wedge\theta^{A_{2}}\wedge...\wedge\theta^{A_{k-1}}\,\right)(\,W_{1},...,\,W_{k-1})\,.$ (62)
The fact that it has to be zero for all the $W_{j}$ gives
$X_{a_{1}\,A_{2}\,...\,A_{k-1}}\,=\,0\,.$ (63)
The same argument can be iterated by considering all the other possible choices for the $W_{j}$. One ends up with
[…]
proving that $X$ is tangent to $\mathcal{M}$ and, thus, that $\mathcal{M}$ is a (k1)(k-1)-coisotropic submanifold of a ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$. ∎
@@
Note that, unlike the classical case $k=2$ [Got82], the form $\tilde{\omega}$ remains non-degenerate throughout the entire ${\Lambda^{k-1}}^{\perp}_{R}(\mathcal{M})$. Consequently, there is no need to restrict to a tubular neighborhood of the zero-section of $\tau$ (as in [Got82]) or to impose additional assumptions such as the flatness of $P$ (see [DGMS93, CDI+22b]) to ensure the non-degeneracy of $\tilde{\omega}$.
Remark 2.3.
Another key difference between the case $k>2$ analyzed here and the classical case $k=2$ concerns the uniqueness of the multisymplectic thickening. Indeed, in [Got82], the symplectic thickening is not only guaranteed to exist but also proven to be unique (up to symplectomorphism). In contrast, in the multisymplectic case, uniqueness generally cannot be established, as the following simple counterexample shows.
Consider the pre-$3$-plectic manifold $(\mathbb{R}^{4},\,\omega)$ with coordinates
$\left\{\,x^{1},\,x^{2},\,x^{3},\,x^{4}\,\right\}\,$, (65)
where
$\omega\,=\,{\rm d}x^{1}\wedge{\rm d}x^{2}\wedge{\rm d}x^{3}\,$, (66)
with kernel given by
$\mathrm{ker}\,\omega\,=\,\langle\,\left\{\,\frac{\partial}{\partial x^{4}}\,\right\}\,\rangle\,.$ (67)
@@
$\left\{\,x^{1},\,x^{2},\,x^{3},\,x^{4},\,x^{5}\,\right\}\,$, (68)
where
$\tilde{\omega}\,=\,{\rm d}x^{1}\wedge{\rm d}x^{2}\wedge{\rm d}x^{3}+{\rm d}x^{1}\wedge{\rm d}x^{4}\wedge{\rm d}x^{5}\,$, (69)
@@
$\left\{\,x^{1},\,x^{2},\,x^{3},\,x^{4},\,x^{5},\,x^{6}\,\right\}\,$, (70)
where
$\tilde{\omega}\,=\,{\rm d}x^{1}\wedge{\rm d}x^{2}\wedge{\rm d}x^{3}+{\rm d}x^{1}\wedge{\rm d}x^{4}\wedge{\rm d}x^{5}+{\rm d}x^{2}\wedge{\rm d}x^{4}\wedge{\rm d}x^{6}\,.$ (71)
They are both $3$-plectic manifolds, and one can easily check that the $(\mathbb{R}^{4},\,\omega)$ is a $2$-coisotropic submanifold of both of them.
Therefore, one concludes that establishing an existence result—depending only on the (immaterial) choice of a complement, as in Gotay’s construction—is the most canonical result one can achieve in the multisymplectic context, where uniqueness is not guaranteed.
3 Application: A Scalar Field Theory with Anisotropic Mass
As an application of Theorem˜2.1, we consider a scalar field theory on a 2-dimensional space-time in the multisymplectic Hamiltonian framework. We refer to [CDI+24] and references therein for the theory behind the multisymplectic approach to Hamiltonian field theories.
In the example we are considering, the space-time is the $2$-dimensional manifold $\mathbb{R}^{2}$ with coordinates $\left\{\,x,\,t\,\right\}$. The multisymplectic phase space is the reduced dual of the first-order jet bundle of the trivial bundle $\pi\colon E\,=\,\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{R}^{2}$. It is a $5$-dimensional manifold $\mathcal{M}$ isomorphic to $\mathbb{R}^{5}$ with coordinates $\left\{\,x,t,u,\rho^{x},\rho^{t}\right\}$, where $u$ represents the scalar field coordinate and $\rho^{x}$, $\rho^{t}$ are the so-called multimomenta coordinates. The fields of the theory are the sections of the bundle $\mathcal{M}\to\mathbb{R}^{2}$, say
$\chi\;:\;\;(x,\,t)\mapsto(x,\,t,\,\phi(x,t),\,P^{x}(x,t),\,P^{t}(x,t))\,.$ (72)
Instead of the multisymplectic form
$\omega\,=\,{\rm d}\rho^{t}\wedge{\rm d}u\wedge{\rm d}x+{\rm d}\rho^{x}\wedge{\rm d}u\wedge{\rm d}t-\rho^{t}{\rm d}\rho^{t}\wedge{\rm d}x\wedge{\rm d}t-\rho^{x}{\rm d}\rho^{x}\wedge{\rm d}x\wedge{\rm d}t\,$, (73)
corresponding to the Hamiltonian
$H\,=\,\frac{1}{2}\left(\,(\rho^{x})^{2}+(\rho^{t})^{2}\,\right)\,$, (74)
we will consider the pre-multisymplectic structure
$\omega={\rm d}\rho^{x}\wedge{\rm d}u\wedge{\rm d}t-\rho^{x}{\rm d}\rho^{x}\wedge{\rm d}x\wedge{\rm d}t\,$, (75)
corresponding to a free scalar field in a singular limit where its mass becomes infinite in the time-like direction.
The form $\omega$ is evidently closed, but it is degenerate. Its kernel is easily seen to be
$K\,=\,\langle\left\{\,\frac{\partial}{\partial x}+\rho^{x}\frac{\partial}{\partial u},\,\frac{\partial}{\partial\rho^{t}}\,\right\}\rangle\,.$ (76)
We now apply Theorem˜2.1 for $k=3$. In this simple example, we will consider the connection $P$ to provide the following complement for $K$
$H\,=\,\langle\left\{\,\frac{\partial}{\partial t},\,\frac{\partial}{\partial u},\,\frac{\partial}{\partial\rho^{x}}\,\right\}\rangle\,.$ (77)
The multisymplectic thickening is the manifold $\tilde{\mathcal{M}}={\Lambda^{2}}^{\perp}_{R}(M)$, i.e., the bundle of transversal 2-forms w.r.t. $R\,=\,\mathbb{1}-P$.
For the connection chosen, the $1$-forms $\theta^{A}$ of Section˜1 and Section˜2 read $\{dx,d\rho^{t}\}$, namely
$P\,=\,\left(\,\frac{\partial}{\partial x}+\rho^{x}\frac{\partial}{\partial u}\,\right)\otimes{\rm d}x+\frac{\partial}{\partial\rho^{t}}\otimes{\rm d}\rho^{t}\,.$ (78)
A transversal $2$-form is a $2$-form containing at least one $\theta^{A}$. This introduces $7$ auxiliary fields, which are the coordinates on the fibers of $\tilde{\mathcal{M}}$
$\left\{\,p_{xt},\,p_{xu},\,p_{x\rho^{x}},\,p_{x\rho^{t}},\,p_{\rho^{t}t},\,p_{\rho^{t}u},\,p_{\rho^{t}\rho^{x}}\,\right\}\,.$ (79)
The manifold $\tilde{\mathcal{M}}$ thus reads $\mathbb{R}^{12}$ with coordinates
$\left\{\,x,\,t,\,u,\,\rho^{x},\,\rho^{t},\,p_{xt},\,p_{xu},\,p_{x\rho^{x}},\,p_{x\rho^{t}},\,p_{\rho^{t}t},\,p_{\rho^{t}u},\,p_{\rho^{t}\rho^{x}}\,\right\}\,.$ (80)
The new, non-degenerate $3$-form on $\tilde{\mathcal{M}}$ is $\tilde{\omega}=\tau^{*}\omega+d\Theta_{0}$, where $\Theta_{0}$ is the tautological 2-form:
$\begin{split}\Theta_{0}=&p_{xt}{\rm d}x\wedge{\rm d}t+p_{xu}{\rm d}x\wedge{\rm d}u+p_{x\rho^{x}}{\rm d}x\wedge{\rm d}\rho^{x}+p_{x\rho^{t}}{\rm d}x\wedge{\rm d}\rho^{t}+\\ +&p_{\rho^{t}t}{\rm d}\rho^{t}\wedge{\rm d}t+p_{\rho^{t}u}{\rm d}\rho^{t}\wedge{\rm d}u+p_{\rho^{t}\rho^{x}}{\rm d}\rho^{t}\wedge{\rm d}\rho^{x}\,.\end{split}$ (81)
3.1 Equations of motion
Here, we will see the role of the regularization procedure in providing a fully determined system of equations of motion out of an implicit underdetermined one, showing that the regularization has the meaning of fixing the undetermined gauge in the equations of motion.
The equations of motion on $(\mathcal{M},\,\omega)$ are derived in the multisymplectic approach [CDI+24] from the condition $\chi^{\star}(i_{V}\omega)=0$ for any vertical vector field on the fibration $\mathcal{M}\to\mathbb{R}^{2}$
$V\,=\,V_{u}\frac{\partial}{\partial u}+V_{\rho^{x}}\frac{\partial}{\partial\rho^{x}}+V_{\rho^{t}}\frac{\partial}{\partial\rho^{t}}\,.$ (82)
A straightforward calculation shows that equations of motion read
[…]
As expected, since the system is degenerate, some degrees of freedom, the momentum $P^{t}$ in particular, are completely undetermined, namely, the system exhibits what is called a gauge symmetry. As a consequence, the field $\phi$ obeys
$\frac{\partial^{2}\phi}{\partial x^{2}}\,=\,0\,$, (84)
but its evolution along the $t$-coordinate is completely free.
[…]
$\tilde{\chi}\;:\;\;(x,\,t)\mapsto(x,\,t,\,\phi(x,t),\,P^{x}(x,t),\,P^{t}(x,t),\Pi_{xt}(x,t),\,\Pi_{xu}(x,t),...,\,\Pi_{\rho^{t}\rho^{x}}(x,t))\,$, (85)
and for any vertical vector field $\tilde{V}$.
Remark 3.1.
We are well aware that a rigorous application of this procedure raises a fundamental structural question: does the multisymplectic thickening $\tilde{\mathcal{M}}$ of a multisymplectic phase space $\mathcal{M}$ (which is, by construction, the reduced dual of a jet bundle) inherit a similar structure? In the corresponding Lagrangian formalism, this is equivalent to asking under which conditions the thickening of a jet bundle is itself a jet bundle.
A thorough geometric analysis of this problem is required to ensure the global consistency of the regularized field theory. However, such an investigation lies beyond the scope of this paper, whose main goal is to establish the coisotropic embedding theorem itself. This structural problem will be addressed in a future work. For the purely illustrative purpose of this example, we proceed by assuming that a consistent description of the regularized fields as sections $\tilde{\chi}$ of a suitable bundle $\tilde{\mathcal{M}}\to\mathbb{R}^{2}$ exists.
A long but straightforward computation shows that in this case, the equations of motion read
[…]
that amount to
[…]
where now, the dynamics of $\phi$ along the direction $t$ has been fixed to be constant.
Additionally, from $\tilde{\chi}^{\star}(i_{\tilde{V}}\tilde{\omega})\,=\,0$ one gets a set of "dynamical" equations for the additional, un-physical fields $\Pi_{xt},...,\,\Pi_{\rho^{t}\rho^{x}}$, that we avoid to write explicitely since they will be neglected.
This example illustrates how the coisotropic embedding procedure takes a pre-multisymplectic system with a non-trivial kernel, and, thus, with a gauge symmetry, and extends it to a larger, non-degenerate multisymplectic system where all the gauge freedoms are fixed.